Derive an expression for the partition function of an ideal monoatomic gas. Hence obtain an expression for its entropy.
1
Expert's answer
2020-06-12T11:18:57-0400

The dispersion ratio of ultrarelativistic particles has the form



partition function of a monatomic classical gas



Using standard thermodynamic relations, we obtain equations of state


"c_v=3, p=\frac{\theta}{v}"


"c_p=c_v+1=4"

Due to the thermal equation of state of an ideal gas p = θ / v and cv = const, the equation of the adiabatic process has the form


"pv^\gamma=const"

where

"\gamma=\frac{c_p}{c_v}=\frac{4}{3}"